OpenStudy (anonymous):

The expression (-x+5)^2 is equivalent to what?

OpenStudy (goldphenoix):

(-x+5)^2 = (-x+5)(-x+5) To simplify it, you want to do the FOIL method. Do you know how to do the foil method?

OpenStudy (anonymous):

yeah its when you go (-x+5)(x+5) or do a little box

OpenStudy (goldphenoix):

Alright... I'll do this step-by-ste. Fill in the blank. -x * -x = __ -x * 5 = __ 5 * -x = __ 5 * 5 = __

OpenStudy (anonymous):

1x^2-5x-5x+25? x^2-10x+25

OpenStudy (anonymous):

@GoldPhenoix ?

OpenStudy (goldphenoix):

Nope. Did you fill in the blank? What does -x * -x give you?

OpenStudy (anonymous):

thats X^2! so the answer would be x^2-10x+25 right

OpenStudy (goldphenoix):

Opps my bad. What's x * 5?

OpenStudy (anonymous):

5x

OpenStudy (goldphenoix):

Great. 5x + (-5x) = ?

OpenStudy (goldphenoix):

Omg. Nevermind. It is x^2-10x+25. I got confused when you said (-x+5)(x+5)

OpenStudy (anonymous):

so i was right! :D

OpenStudy (goldphenoix):

Yep. :)